OpenStudy (anonymous):

Hey can some one help me its a easy question but i seem to not get something so i have to find value of t. [2 / (t-3) ] + [1 / (t+4)] =0 2t+8+t-3 / (t-3) (t-4)=0

OpenStudy (amistre64):

is this a system of equations? or is this part of your own attempt?

OpenStudy (anonymous):

Its is my attempt the second line the first line is question and we have to find t value

OpenStudy (amistre64):

[2 / (t-3) ] + [1 / (t+4)] =0 multiply thru by say (t+4) [2(t+4) / (t-3) ] + 1 =0 multiply thru by (t-3) 2(t+4) + 1(t-3) =0

OpenStudy (amistre64):

distrbute, and collect your like terms

OpenStudy (amistre64):

in your attempt, realize that a fraction is equal to 0 if the top is equal to 0 ... and the bottom isnt 0 0/n = 0 for some nonzero n

OpenStudy (anonymous):

I didnt get it where u got + 1 from in third line

OpenStudy (amistre64):

[2 / (t-3) ] + [1 / (t+4)] =0 multiply thru by say (t+4) <------ [2 (t+4)/ (t-3) ] + [1 (t+4)/ (t+4)] =0(t+4) [2 (t+4)/ (t-3) ] + 1 =0

OpenStudy (amistre64):

since t+4 is not zero .... its on the bottom of a fraction to start with ... we can multiply thru and cancel it out

OpenStudy (amistre64):

if we use your method: 2t+8+t-3 / (t-3) (t-4)=0 only if the top part is 0 soo 2t+8+t-3 = 0